the hypotenuse of a right triangle is 26 cm. the sum of the two other sides is 34 cm. find the lengths of the other two sides of the triangle. how do i solve this ?
Answer by stanbon(75887)

the hypotenuse of a right triangle is 26 cm. the sum of the two other sides is 34 cm. find the lengths of the other two sides of the triangle.
---------------------------------
let one side be x
The other side is 34-x
-----------------------------
Pythagoras:
x^2 + (34-x)^2 = 26^2
x^2 + 1156 - 68x + x^2 = 676
2x^2 -68x + 480 = 0
x^2 - 34x + 240 = 0
(x-24)(x-10) = 0
x = 24 or x = 10
------------------------------
One side is 10 cm the other is 24 cm.
=========================================
Cheers,
stan H.